# RSNet: A Light Framework for The Detection of SAR Ship Detection

### v5

@@ I Introduction
With SAR image data scale and complexity increasing, traditional image processing methods are becoming insufficient [9]. Recent advances in convolutional neural networks (CNNs) and Transformer architectures have significantly improved target detection [10], [11]. These deep learning methods are generally categorized into candidate region-based and regression-based approaches, with the latter excelling in real-time detection by directly regressing targets at multiple scales. For SAR object detection, one-stage methods strike a promising balance between speed and accuracy; however, due to the computational and energy limitations of airborne platforms, a lightweight model framework is essential for practical deployment.
@@ II-D LS Detector
Figure 4: The heat-map visualization results feature three rows: the first row illustrates the ground truth, the second row presents the baseline (YOLOv8n), and the third row highlights our proposed method. Red circles represent missed detections.
@@
× × 91.2 66.5 2.43 7.4
× 91.1 65.8 2.13 6.8
91.2 67.6 1.49 5.1
@@ III-A Datasets
To evaluate the performance of our model, we utilized two publicly available SAR remote sensing datasets: SSDD [19] and HRSID [20]. The SSDD consists of 1160 images, with 928 allocated for training and 232 for testing, as outlined in the original publication. In the case of the HRSID, we adhered to the official partitioning, employing 3642 images for training and 1962 for testing purposes.
@@ III-B Training Details
RetinaNet 78.9 53.6 36.33 128
YOLOv6n 88.2 62.8 4.23 11.8
YOLOv7-tiny 85.4 57.2 6.01 13.2
@@
RetinaNet 90.2 60.0 36.33 128
YOLOv6n 96.9 71.2 4.23 11.8
YOLOv7-tiny 96.4 66.5 6.01 13.2
@@ III-C Evaluation Metrics
To evaluate the proposed model’s performance, we used mean average precision at IoU = 0.5 ($\mathrm{mAP_{.50}}$) and IoU = 0.5:0.05:0.95 ($\mathrm{mAP_{.50:.95}}$), along with the number of parameters (Params) and floating-point operations per second (FLOPs). The mAP metrics assess detection accuracy, while Params and FLOPs indicate parameters and computational complexity.
@@ III-D Ablation Experiments
To validate the module’s effectiveness, we conducted ablation experiments on the HRSID dataset, as shown in Table II. We also performed feature visualization experiments, illustrated in Fig. 4, which demonstrate that our model effectively identifies real targets while reducing the influence of surrounding noise.

### v7

@@ I Introduction
With SAR image data scale and complexity increasing, traditional image processing methods are becoming insufficient [9]. Recent advances in convolutional neural networks (CNNs) and Transformer architectures have significantly improved target detection [10], [11], [12], [13]. These deep learning methods are generally categorized into candidate region-based and regression-based approaches, with the latter excelling in real-time detection by directly regressing targets at multiple scales. For SAR object detection, one-stage methods strike a promising balance between speed and accuracy; however, due to the computational and energy limitations of airborne platforms, a lightweight model framework is essential for practical deployment.
@@ II-D LS Detector
Figure 4: The heat-map visualization results feature three rows: the first row illustrates the ground truth, the second row presents the baseline1 (YOLOv8n), and the third row highlights our proposed method. Red circles represent missed detections.
@@
Figure 5: Performance Comparison of Three Baseline Models and the RSNet-Enhanced Model in Terms of Preprocessing Time and Inference Time.
@@
× × 91.2 (0.6 $\uparrow$) 66.5 (0.4 $\downarrow$) 2.43 (0.58 $\downarrow$) 7.4 (0.7 $\downarrow$)
× 91.1 65.8 (1.1 $\downarrow$) 2.13 (0.88 $\downarrow$) 6.8 (1.3 $\downarrow$)
91.2 (0.1 $\uparrow$) 67.6 (0.7 $\uparrow$) 1.49 (1.52 $\downarrow$) 5.1 (3.0 $\uparrow$)
Baseline2 (YOLOv9n)
× × × 90.5 66.1 2.01 7.8
× × 91.1 (0.6 $\uparrow$) 66.9 (0.8 $\uparrow$) 1.65 (0.36 $\downarrow$) 7.2 (0.6 $\downarrow$)
× 90.3 (0.2 $\downarrow$) 65.6 (0.5 $\downarrow$) 1.24 (0.77 $\downarrow$) 5.9 (1.9 $\downarrow$)
91.1 (0.6 $\uparrow$) 66.6 (0.5 $\uparrow$) 0.77 (1.24 $\downarrow$) 4.5 (3.3 $\downarrow$)
Baseline3 (YOLOv11n)
× × × 89.7 65.8 2.59 6.4
× × 90.0 (0.3 $\uparrow$) 66.2 (0.4 $\uparrow$) 2.01 (0.58 $\downarrow$) 5.2 (1.2 $\downarrow$)
× 90.0 (0.3 $\uparrow$) 66.1 (0.3 $\uparrow$) 1.30 (1.29 $\downarrow$) 5.1 (1.3 $\downarrow$)
91.1 (0.4 $\uparrow$) 67.0 (1.2 $\uparrow$) 1.05 (1.54 $\downarrow$) 4.9 (1.5 $\downarrow$)
@@ III-A Datasets
We evaluated our model on two public SAR datasets: SSDD [21] and HRSID [22]. SSDD includes 1160 images, with 928 for training and 232 for testing as per the original split. HRSID follows the official split, using 3642 training and 1962 testing images.
@@ III-B Training Details
RetinaNet [20] 78.9 53.6 36.33 128
YOLOv6n [23] 88.2 62.8 4.23 11.8
YOLOv7-tiny [24] 85.4 57.2 6.01 13.2
@@
RetinaNet [20] 90.2 60.0 36.33 128
YOLOv6n [23] 96.9 71.2 4.23 11.8
YOLOv7-tiny [24] 96.4 66.5 6.01 13.2
@@ III-C Evaluation Metrics
Model performance was assessed using $\mathrm{mAP_{.50}}$ and $\mathrm{mAP_{.50:.95}}$ for detection accuracy, alongside parameter count (Params) and FLOPs to quantify model complexity and computational efficiency.
@@ III-D Ablation Experiments
To validate the module’s effectiveness, we conducted ablation experiments on the HRSID dataset based on YOLOv8n, YOLOv9n and YOLOv11n, as shown in Table II. We also performed feature visualization experiments, illustrated in Fig. 4, which demonstrate that our model effectively identifies real targets while reducing the influence of surrounding noise. To evaluate the performance of the proposed framework under computational constraints, we test the baseline models and their RSNet-enhanced models on an Intel i5-8265U CPU (TDP = 15W) at 20°C under CPU-only inference for fair comparison. Results are shown in Fig. 5.